# Distinguishing Predictive and Generative AI in Regulation

### v2

@@ 3. What makes generative AI distinctive?
Even when tasks are well-specified, the real-world performance and impact of generative models are shaped by how those tasks are carried out via fine-tuning, prompting, and other forms of downstream adaptations (Raji et al., 2021; Liao and Xiao, 2023; Bowman and Dahl, 2021; Chi et al., 2025). These post-training modifications are often highly variable and unpredictable (Betley et al., 2025), leading the same model to exhibit markedly different behavior even when used to perform what seem to be the same or similar tasks. Given how sensitive evaluation results can be to these downstream choices, it can be exceedingly difficult to reliably predict performance based on evaluations that only consider a limited set of possible choices.
@@
Evaluation at the foundation model level falls prey to what Selbst et al. calls the ”framing trap”—the failure to account for the broader sociotechnical system in which AI operates (Selbst et al., 2019). With generative AI, the framing trap is largely unavoidable if one seeks to evaluate the foundation model. The sociotechnical context of the AI system varies with each use case, shaped by the distinct patterns of user interactions, institutional norms, and task-specific objectives (Blodgett et al., 2020). As a result, even if it is feasible to develop robust, standardized evaluation methods for the foundation model itself, evaluating foundation models in isolation from the context of its uses cannot adequately surface concrete harms or support meaningful mechanisms of accountability.
@@
While past regulatory requirements around AI audits and impact assessments have hinged on the assumption that developers can meaningfully evaluate AI systems prior to deployment, generative AI challenges the reliability of current evaluation methods as regulatory tools. One example that highlights its limitations is California’s S.B. 1047. The vetoed state bill would have required model developers “to assess whether the covered model is reasonably capable of causing or materially enabling critical harm” before making it publicly available (California Senate, 2024). The problem, however, is that evaluating harm at the model level, without accounting for downstream use cases, fails to consider the role of user interactions, implementation decisions, and the organizational processes in shaping the real-world impacts of model use. As such, regulatory proposals hitched to pre-deployment evaluations of foundation models may misdiagnose where and how harm actually materialized, leading to ineffective or misguided oversight.
@@
3.3. Generative AI opens up new areas of legal concern that were not implicated by predictive AI.
@@
Copyright offers the most prominent example to the general public. While predictive AI can raise copyright concerns at the input stage—whether training on protected works constitutes infringement (Levendowski, 2018; Sobel, 2017)—this was not the primary legal concern with the technology. Generative AI not only magnifies concerns about infringement in training, but also implicates the reproduction right due to its ability to output text, images, and videos, which predictive AI did not (Lee et al., 2024; Henderson et al., 2023; Lemley, 2023). AI companies currently face major infringement lawsuits based on using protected content in model training and producing outputs that closely resemble copyrighted works (Grynbaum and Mac, 2023; Authors Guild et al. v. OpenAI Inc. et al., 2023; Chabon et al. v. OpenAI, Inc., 2023; Getty Images , Inc. v. Stability AI, Inc.(2023), US).
@@
Beyond copyright concerns, generative AI has drawn scrutiny for its role in spreading defamation and misinformation. Generative systems regularly fabricate plausible-sounding falsehoods, producing content with invented names, fake quotations, and non-existent sources (OpenAI, 2023). Already, lawsuits have been filed over alleged libel created by generative AI programs (Mark Walters v. OpenAI, 2024; v. Microsoft Corporation, 2023). As generative systems are increasingly integrated into search engines and virtual assistants, the risk of reputational harm and public deception intensifies. In more extreme cases, generative systems have been used to commit fraud. Synthetic voice technologies have been used to clone speech patterns of real individuals, enabling impersonation through robocalls (Federal Communications Commission, 2024; Barrington et al., 2025). Hyperrealistic deepfake videos and AI-generated images have similarly been used to deceive viewers and distort perception, and pose particular difficulties for evidentiary processes in courts (Nightingale and Farid, 2022; Bohacek and Farid, 2025). The proliferation of synthetic content raises the question of whether and when AI developers should be held liable for content generated by its systems (Volokh, 2023; Lee et al., 2024).
@@
While predictive systems have long played a role in intelligence analytics and surveillance, generative AI introduces new concerns in chemical, biological, radiological, and nuclear (CBRN) security (UK Government, 2024; National Institute of Standards and Technology, 2024). Traditionally, the development of CBRN weapons has required high levels of expertise, access to restricted materials, and coordination across domains. However, generative systems could lower informational barriers and expand the capabilities of malicious actors by synthesizing vast bodies of literature and proposing novel ideas for CBRN weapon development (Soice et al., 2023; Mouton et al., 2024; Gopal et al., 2023). While the actual realization of such threats still depends on real-world capabilities (Peppin et al., 2025), the possibility that generative models could support CBRN threats has raised concerns about their dual-use potential, driving new discussions about risk mitigation, early warning systems, and access controls (OpenAI, 2024; The White House, 2023; Batalis, 2024; National Academies of Sciences, Engineering, and Medicine, 2025).
@@
The legal challenges posed by generative AI in new domains highlight the limitations of treating AI as a singular regulatory target. These issues fall under distinct legal regimes, each with its own doctrines, precedents, and norms that have developed over decades. For example, copyright law has long grappled with questions of authorship, originality, and fair use, and has developed specific doctrinal tests to differentiate between protected creative works and derivative copies (17 U.S.C. § 102(nd), b). CBRN security governance has been grounded in the precautionary principle, the practice of taking preventative action in the face of uncertainty. It has relied on nonproliferation regimes, material control measures, and international oversight to restrict the development and spread of high-risk threats (United Nations, nd). Each of these domains has also drawn on distinct communities of experts and stakeholders with different priorities, risk tolerances, and terminologies. The differences in how each community conceptualizes and addresses the challenges posed by generative AI mean that regulatory approaches that work well in one domain may be ineffective or counterproductive in another. As a result, effective governance for generative AI will require broader and more diverse domain knowledge than is required for predictive AI alone. Attempts to consolidate these distinct legal challenges into a single, overarching AI policy framework risks not only oversimplifying the regulatory landscape but also overlooking the stakeholders and experts best positioned to guide policy decisions.
@@
3.4. Generative AI is built on a highly distributed value chain that complicates accountability.
@@
Predictive AI systems typically exhibit a tightly integrated value chain, from data curation to feature selection to model development. The system’s single purpose and structure are understood at the time of creation, and the creation itself is often done collaboratively between developer and deployer (De Silva and Alahakoon, 2022). Thus, the universe of potential harms is either known, or at least foreseeable, to both the developer and deployer of the system.
@@
Generative AI complicates this model. As model development is no longer tied to specific tasks, model providers focus on building large, incredibly complex, general purpose models that are designed to be customized and reused. Then, in the minimally complex case, a deployer will fine-tune the system for their use case and then deploy it. There are several important implications to this change for legal concepts related to responsibility and accountability.
@@
The first is that the conceptual distance between the model developer’s work and the ultimate harm is much larger. Where in a predictive model, the developer will be aware of the final use, and will often work in concert with their customer, now they are building general purpose technologies, and many ultimate uses will be hard, if not impossible, to foresee (Widder and Nafus, 2022). Additionally, the complexity of these models renders specific outcomes likely unforeseeable, even where the general domain of use can be predicted in advance (Cobbe et al., 2023). Foreseeability, in turn, is a bedrock limitation to legal liability in most instances (Hart and Honoré, 1985). This translates to a general aversion to hold creators of general-purpose systems with substantial lawful uses liable for downstream harms (Sony Corp. of America v. Universal City Studios, Inc., 1984). Telephone creators are not held liable for crimes organized over the phone, nor are computer makers for harms researched and planned on laptops.
@@
The second is that there are more—and more specialized—roles in the overall value chain. This is not inherently difficult for a legal regime to handle. For example, the EU’s AI Act lays out different sets of responsibilities for providers, deployers, importers, and distributors (European Parliament, 2024). But new classes of intermediate actors—model adapters and integrators—are getting involved. Model adapters fine-tune or modify generative models for specific tasks or domains, while integrators incorporate models into applications or systems, shaping how users interact with and experience them. The fluidity of the different value chains and the disaggregation of responsibilities into different roles can make it difficult for a statute to stay current (Engler and Renda, 2022).
@@
This fragmentation of roles is not inherently problematic for a legal regime. In product liability, for instance, all parties in the chain of commerce are held liable for an injury to a plaintiff, and the court lets them sort out who ultimately pays. In the EU AI Act, a deployer who makes substantial modifications or changes the purpose of a deployed system becomes a provider as a legal matter.(European Parliament, 2024, Article 25). As long as someone is liable in the end, then oversight can be accomplished. What’s different here is that lines between actors in the generative AI value chain are often blurry, making the specific questions of how responsibility is divided more difficult, raising the cost of oversight. Sticking with products liability as an example, under the so-called “component parts doctrine,” a supplier of a component in a larger products is liable for the harm only if the component itself is defective, the integration causes the component to be defective, or the maker of the component substantially participates in the integration (of Torts: Product Liability, 1998) This doctrine relies on a continued conceptual separation between the component and the larger product. What’s different now is that the downstream uses actually morph the product at issue. Once another actor has adapted it for some purpose, there is no way to once again separate the new version of the AI into its components, to understand who is responsible for what ultimate harm. This will lead to costly litigation.
@@
Third, the extra roles and distributed global nature of the AI value chain may pose jurisdictional difficulties in legal accountability. Models can be developed in one country and be ported across jurisdictions trivially. Yet the legality of their training or deployment may differ by jurisdiction. For example, if personal data of any EU residents was used to train foundation models, the training likely violated the GDPR, so what should EU regulators make of someone who independently fine-tunes one of the existing models for a local purpose? These challenges are amplified in the case of open-source or open-weight models, which are published and freely accessible anywhere. In contrast, these kinds of concerns are less common with predictive models, as they are typically bespoke systems trained on proprietary data and developed with infrastructure confined to a single jurisdiction.
None of this means that upstream providers should be exempt from responsibility for downstream harms. For one, they maintain access to core system components, and retain the ability to control the systems they release. In addition, because there are only a small number of upstream providers, they have immense market power (UK Competition and Markets Authority, 2023).
@@
Generative AI has an unprecedented demand for compute and data. While earlier predictive systems might be trained on thousands or millions of examples, generative models rely on billions or trillions of data points (AI Now Institute, 2023). On average, training large-scale generative models requires roughly 100 times more compute than earlier systems (Sevilla et al., 2022). As a result, upstream providers have become systematically important within the generative AI value chain, as downstream actors increasingly rely on a few providers for access to pretrained models, large-scale data, and advanced compute infrastructure—resources foundational to building and scaling generative systems (Cobbe et al., 2023). These intensified dependencies give upstream providers an outsized influence over who can build, scale, and commercialize generative AI, a dynamic that was far less pronounced in the context of predictive AI.
@@ 4. Recommendations
Drawing on past lessons from governing predictive AI, we propose three recommendations for policymakers to consider when it comes to generative AI. First, regulators should identify regulatory targets based on the function and consequences of AI systems rather than technical implementation. Second, only in narrow circumstances should the technical properties of AI systems be employed to determine the level of regulatory scrutiny. Third, regulators should examine the entire risk chain of AI applications and strengthen structural constraints that could serve as more effective barriers to misuse.
@@
The challenge of regulating AI exemplifies many of the core sources of contention and difficulties in tech policy. Central to this discourse is the distinction between tech-neutral and tech-specific regulatory approaches. Tech-neutral regulation governs activities and consequences rather than specific technical implementations.
Past legal scholars have advocated for tech-neutral regulation for several reasons (Koops et al., 2006; Pikowsky, 2003; McEniery, 2010; Goodman and Chen, 2011; Kerr, 2010). First, it prevents legal gaps that arise when tech-specific regulation fails to keep pace with technical development (Eichensehr, 2015; McEniery, 2010). With rapid technological advancements, restrictions based on current specifications may become obsolete by the time laws are enacted. Second, it avoids overly broad or narrow definitions that may inadvertently include or exclude technologies, resulting in potentially ineffective or even harmful regulation (Casey and Lemley, 2020; Koops et al., 2006). Tech-neutral regulation instead defines its regulatory scope based on the applications and outcomes of the technology, allowing regulators to target the societal impact of a system in a way that ensures proportionate oversight. Third, tech-neutral regulation is resistant to regulatory arbitrage, preventing actors from evading oversight by switching to alternative technical approaches that achieve the same concerning outcomes (Crootof and Ard, 2021; Birnhack, 2012). Even with well-defined regulatory targets for tech-specific regulation, actors can easily develop functionally similar systems that use different techniques to circumvent the rules.
The case for tech-neutral regulation has taken on new significance in the context of AI. In predictive AI, model development is oriented to specific use cases — this task specificity facilitates compliance with tech-neutral regulation, as we argue in Section 3.1. However, for generative AI, this specificity is lost due to the separation between development and deployment. Instead, the general purpose abilities of generative AI have led to calls for regulation at the level of the model itself, leading to shallow approaches based on model parameters and other technical proxies. We refer to technical proxies for AI systems as measurable characteristics of their inputs and outputs along two axes defined by Bommasani (Bommasani, 2023): (1) resource expenditure (i.e., the amount of compute in FLOPs and the amount of data in modality-specific units), and (2) model properties (i.e. capabilities benchmarks and safety benchmarks).
@@
AI regulation that relies on technical proxies risks not just becoming outdated but being both under- and over-inclusive (Crootof and Ard, 2021). [^5] Considered more precisely, this is a point about legal rules and standards, not tech neutrality (Crootof and Ard, 2021, p. 406). Where regulations are more rule-like, they can be over- and under-inclusive. We make the point here about tech-specificity because we think it is likely that tech-specificity leads to rules more often than standards. This is both because the tendency for tech-specific regulation is already motivated by a desire for easier adminsitrability and because technical standards and specifications are often stated in quantitative terms. Examples of tech-specific standard-based regulation may become more prevalent if governments, for example, mandate that LLMs have output filters that accomplish some goal like fairness or a more complicated set of goals in the vein of Anthropic’s attempt at “Constitutional AI”, (Bai et al., 2022) but for now, we think the assumption holds. Regulation is underinclusive when it leaves out systems that achieve similar concerning outcomes through alternative methods. For instance, relying on compute thresholds to identify frontier AI models may inadvertently exclude systems that cause harm through less resource-intensive methods (e.g., misinformation bots) (Newsom, 2024). Regulation is overinclusive when it captures more systems than intended. For example, the same compute thresholds could sweep in benign applications like climate modeling that happen to share high computational demands. The challenge in defining the regulatory scope is compounded by the broad range of contexts in which the same AI systems can be deployed to serve vastly different purposes (Harris, 2021). The generality and adaptability of generative AI therefore underscore the need for regulatory frameworks that account for the specific contexts and outcomes of their use (Murdick et al., 2020).
@@
One example that highlights the limits of tech-specific regulation is the European Union’s legislation on genetically modified organisms (GMO) (Parliament and the Council of the European Union, 2001). The legislation targets crops that have been altered by recombinant DNA technology to address public concerns over potential health risks. In response, seed producers have turned away from recombinant DNA technology toward mutagenesis, using mutagenic chemicals or radiation treatments to induce random genetic changes in plants (Kaskey, 2013a). While these methods are not considered recombinant DNA technology, they achieve the same outcome of producing genetically modified crops with commercially desirable traits. Ironically, mutagenic crops may pose comparable or even greater risks than GMO crops on food safety and the environment (Kaskey, 2013b; Morris, 2007). This case demonstrates how tech-specific regulation can incentivize the development of what Dan Burk calls “perverse innovations” that “ingeniously dodge the intended outcome of regulation, while still formally adhering to the text of the regulation (Burk, 2016).”
4.2. Identifying regulatory targets by technical specifications can make sense as a last resort.
In recent years, the AI regulatory landscape has seen a chorus of concerns over AI safety and security risks (Tobin, 2024; Lovely, 2024; Tufekci, 2025; Future of Life Institute, 2023). These concerns have led to a surge in tech-specific proposals that aim to regulate AI systems, with model properties as technical proxies to their capacity to demonstrate greater impact or precipitate harm in society (European Parliament, 2024; U.S. House Committee on Foreign Affairs, 2024; The White House, 2023; California Senate, 2024).
@@
However, tech-specific regulation is not without its limitations. As we highlighted in the previous section, its application should be carefully constrained to avoid unintended consequences, such as overregulating benign use cases and creating regulatory loopholes. Therefore we argue that technical proxies should be employed for regulation only in narrow circumstances where three conditions are met: (1) tech-neutral regulation is demonstrably impractical in addressing the risk of concern, (2) the risk being addressed is severe, plausible, and attributable to the existence and configuration of the technology, and (3) the benefits from tech-specific regulation outweigh the potential harms. We now expand on each of these three conditions.

### v3

@@ 3. What makes general-purpose AI distinctive?
Even when tasks are well-specified, the real-world performance and impact of GPAI models are shaped by how those tasks are carried out via fine-tuning, prompting, and other forms of downstream adaptations (Raji et al., 2021; Liao and Xiao, 2023; Bowman and Dahl, 2021; Chi et al., 2025). These post-training modifications are often highly variable and unpredictable (Betley et al., 2025), leading the same model to exhibit markedly different behavior even when used to perform what seem to be the same or similar tasks. Given how sensitive evaluation results can be to these downstream choices, it can be exceedingly difficult to reliably predict performance based on evaluations that only consider a limited set of possible choices.
@@
Evaluation at the foundation model level falls prey to what Selbst et al. call the ”framing trap”—the failure to account for the broader sociotechnical system in which AI operates (Selbst et al., 2019). With GPAI, the framing trap is largely unavoidable if one seeks to evaluate the foundation model. The sociotechnical context of the AI system varies with each use case, shaped by the distinct patterns of user interactions, institutional norms, and task-specific objectives (Blodgett et al., 2020). As a result, even if it is feasible to develop robust, standardized evaluation methods for the foundation model itself, evaluating models in isolation from the context of its uses cannot adequately surface concrete harms or support meaningful mechanisms of accountability, which would be seen or apply only to the combination of model and application context.
@@
While past regulatory requirements around AI audits and impact assessments have hinged on the assumption that developers can meaningfully evaluate AI systems prior to deployment, GPAI challenges the reliability of current evaluation methods as regulatory tools. One example that highlights its limitations is California’s S.B. 1047. The vetoed state bill would have required model developers “to assess whether the covered model is reasonably capable of causing or materially enabling critical harm” before making it publicly available (California Senate, 2024). The problem, however, is that evaluating harm at the model level, without accounting for downstream use cases, fails to consider the role of user interactions, implementation decisions, and the organizational processes in shaping the real-world impacts of model use. As such, regulatory proposals hitched to pre-deployment evaluations of GPAI models may misdiagnose where and how harms actually materialize, leading to ineffective or misguided oversight.
@@
3.3. General-purpose AI opens up new areas of legal concern that were not implicated by task-specific AI.
@@
Recent developments make clear that GPAI poses legal challenges largely absent from task-specific AI. As such, the responses of existing legal regimes will necessarily differ, each requiring distinct domain-specific expertise. Copyright offers the most prominent example to the general public. While task-specific AI can raise copyright concerns at the input stage—whether training on protected works constitutes infringement (Levendowski, 2018; Sobel, 2017)—this was not the primary legal concern with the technology. GPAI not only magnifies concerns about infringement in training, but also implicates the reproduction right due to its ability to output text, images, and videos, which task-specific AI did not (Lee et al., 2024; Henderson et al., 2023; Lemley, 2023). AI companies currently face major infringement lawsuits based on using protected content in model training and producing outputs that closely resemble copyrighted works (Grynbaum and Mac, 2023; Authors Guild et al. v. OpenAI Inc. et al., 2023; Chabon et al. v. OpenAI, Inc., 2023; Getty Images (US), Inc. v. Stability AI, Inc., 2023).
@@
Beyond copyright concerns, GPAI has drawn scrutiny for its role in spreading defamation and misinformation. GPAI systems regularly fabricate plausible-sounding falsehoods, producing content with invented names, fake quotations, and non-existent sources (OpenAI, 2024b). Already, lawsuits have been filed over alleged libel created by GPAI programs (Mark Walters v. OpenAI, 2024; v. Microsoft Corporation, 2023). As GPAI systems are increasingly integrated into search engines and virtual assistants, the risk of reputational harm and public deception intensifies. In more extreme cases, GPAI systems have been used to commit fraud. Synthetic voice technologies have been used to clone speech patterns of real individuals, enabling impersonation through robocalls (Federal Communications Commission, 2024; Barrington et al., 2025). Hyperrealistic deepfake videos and AI-generated images have similarly been used to deceive viewers and distort perception, and pose particular difficulties for evidentiary processes in courts (Nightingale and Farid, 2022; Bohacek and Farid, 2025). The proliferation of synthetic content raises the question of whether and when AI developers should be held liable for content generated by its systems (Volokh, 2023; Lee et al., 2024).
@@
While task-specific systems have long played a role in intelligence analytics and surveillance, GPAI introduces new concerns in chemical, biological, radiological, and nuclear (CBRN) security (UK Government, 2024; National Institute of Standards and Technology, 2024). Traditionally, the development of CBRN weapons has required high levels of expertise, access to restricted materials, and coordination across domains. However, GPAI systems could lower informational barriers and expand the capabilities of malicious actors by synthesizing vast bodies of literature and proposing novel ideas for CBRN weapon development (Soice et al., 2023; Mouton et al., 2024; Gopal et al., 2023). While the actual realization of such threats still depends on real-world capabilities (Peppin et al., 2025), the possibility that GPAI models could support CBRN threats has raised concerns about their dual-use potential, driving new discussions about risk mitigation, early warning systems, and access controls (OpenAI, 2024a; The White House, 2023; Batalis, 2024; National Academies of Sciences, Engineering, and Medicine, 2025).
@@
The legal challenges posed by GPAI in new domains highlight the limitations of treating AI as a singular regulatory target. These issues fall under distinct legal regimes, each with its own doctrines, precedents, and norms that have developed over decades. For example, copyright law has long grappled with questions of authorship, originality, and fair use, and has developed specific doctrinal tests to differentiate between protected creative works and derivative copies (17 U.S.C. § 102(b), n.d.). CBRN security governance has been grounded in the precautionary principle, the practice of taking preventative action in the face of uncertainty. It has relied on nonproliferation regimes, material control measures, and international oversight to restrict the development and spread of high-risk threats (United Nations, n.d.). Each of these domains has also drawn on distinct communities of experts and stakeholders with different priorities, risk tolerances, and terminologies. The differences in how each community conceptualizes and addresses the challenges posed by GPAI mean that regulatory approaches that work well in one domain may be ineffective or counterproductive in another. As a result, effective governance for GPAI will require broader and more diverse domain knowledge than is required for task-specific AI alone. Attempts to consolidate these distinct legal challenges into a single, overarching AI policy framework risks not only oversimplifying the regulatory landscape but also overlooking the stakeholders and experts best positioned to guide policy decisions.
@@
3.4. General-purpose AI is built on a highly distributed value chain that complicates accountability.
@@
Task-specific AI systems typically exhibit a tightly integrated value chain, from data curation to feature selection to model development. The system’s single purpose and structure are understood at the time of creation, and the creation itself is often done collaboratively between developer and deployer (De Silva and Alahakoon, 2022). Thus, the universe of potential harms is either known, or at least foreseeable, to both the developer and deployer of the system.
@@
Generative AI complicates this picture. As model development is no longer tied to specific tasks, providers focus on building large, incredibly complex, general-purpose models that are designed to be customized and reused. Then, in the minimally complex case, a deployer will fine-tune the system for their use case and then deploy it. There are several important implications to this change for legal concepts related to responsibility and accountability.
@@
The first is that the conceptual distance between the model developer’s work and the ultimate harm is much larger. With a task-specific model, developers will be aware of the final use, and will often work in concert with their customers; now, with a GPAI model, they are building general-purpose technologies, and many ultimate uses will be hard, if not impossible, to foresee (Widder and Nafus, 2022). Additionally, the complexity of these models renders specific outcomes likely unforeseeable, even where the general domain of use can be predicted in advance (Cobbe et al., 2023). Foreseeability, in turn, is a bedrock limitation to legal liability in most instances (Hart and Honoré, 1985). This translates to a general aversion to holding creators of general-purpose systems with substantial lawful uses liable for downstream harms (Sony Corp. of America v. Universal City Studios, Inc., 1984). Telephone creators are not held liable for crimes organized over the phone, nor are computer makers for harms researched and planned on laptops.
@@
The second is that there are more—and more specialized—roles in the overall value chain. This is not inherently difficult for a legal regime to handle. For example, the EU’s AI Act lays out different sets of responsibilities for providers, deployers, importers, and distributors (European Parliament, 2024). But new classes of intermediate actors—model adapters and integrators—are getting involved. Model adapters fine-tune or modify GPAI models for specific tasks or domains, while integrators incorporate models into applications or systems, shaping how users interact with and experience them. The fluidity of the different value chains and the disaggregation of responsibilities into different roles can make it difficult for a statute to stay current (Engler and Renda, 2022).
@@
This fragmentation of roles is not inherently problematic for a legal regime. In product liability, for instance, all parties in the chain of commerce are held liable for an injury to a plaintiff, and the court lets them sort out who ultimately pays. In the EU AI Act, a deployer who makes substantial modifications or changes the purpose of a deployed system becomes a provider as a legal matter (European Parliament, 2024, Article 25). As long as someone is liable in the end, then oversight can be accomplished. What’s different here is that lines between actors in the GPAI value chain are often blurry, making the specific questions of how responsibility is divided more difficult and raising the cost of oversight (Yew, 2024). Sticking with products liability as an example, under the so-called “component parts doctrine,” a supplier of a component in a larger products is liable for the harm only if the component itself is defective, the integration causes the component to be defective, or the maker of the component substantially participates in the integration (of Torts: Product Liability, 1998). This doctrine relies on a continued conceptual separation between the component and the larger product after integration. What’s different now is that the downstream uses actually morph the product at issue. Once another actor fine-tunes a model, adapts it to a new context, or integrates it into a larger application, the resulting system cannot be cleanly disaggregated back into its original components to determine who is responsible for which aspects of its behavior. This lack of separability means that courts will need to examine in detail how the system was modified, what changes those modifications introduced, and how those changes are connected to the harm at issue. Untangling this highly convoluted chain of modifications and actors will raise both the duration and expense of litigation.
@@
Third, the extra roles and distributed global nature of the AI value chain may pose jurisdictional difficulties in legal accountability. Models can be developed in one country and be ported across jurisdictions trivially. Yet the legality of their training or deployment may differ by jurisdiction. For example, if personal data of any EU residents was used to train GPAI models, the training likely violated the GDPR, so what should EU regulators make of someone who independently fine-tunes one of the existing models for a local purpose? Indeed, Yew et al. (Yew et al., 2025) argue that companies are likely to exploit the geographical scoping in the AI Act to continue to provide services to customers in the EU while avoiding governance requirements in the Act. Current US-EU disputes over whether the Digital Services Act should apply to US-based LLM providers, and its implication for free speech protections within the United States are yet another example of the cross-border jurisdictional challenges that GPAI presents. These challenges are amplified in the case of open-source or open-weight models, which are published and freely accessible anywhere. In contrast, these kinds of concerns are less common with task-specific models, as they are typically bespoke systems trained on proprietary data and developed with infrastructure confined to a single jurisdiction.
@@
None of this means that upstream providers should be exempt from responsibility for downstream harms. For one, they maintain access to core system components, and retain the ability to control the systems they release. Because there are only a small number of upstream providers, they have immense market power (UK Competition and Markets Authority, 2023). Further, GPAI has an unprecedented demand for compute and data. While earlier task-specific systems might be trained on thousands or millions of examples, GPAI models rely on billions or trillions of data points (AI Now Institute, 2023). On average, training large-scale GPAI models requires roughly 100 times more compute than earlier systems (Sevilla et al., 2022). As a result, upstream providers have become systematically important within the GPAI value chain, as downstream actors increasingly rely on a few providers for access to pretrained models, large-scale data, and advanced compute infrastructure—resources foundational to building and scaling GPAI systems (Cobbe et al., 2023). These intensified dependencies give upstream providers an outsized influence over who can build, scale, and commercialize GPAI, a dynamic that was far less pronounced in the context of task-specific AI.
@@ 4. Recommendations
Whether emerging technology drives policy has been a longstanding question in technology law, including AI governance (Crootof and Ard, 2021; Calo, 2015; Jones, 2018). Drawing on past lessons from tech policymaking, we propose three recommendations for policymakers to consider when it comes to GPAI. First, regulators should identify regulatory targets based on the function and consequences of AI systems rather than technical implementation. Second, only in narrow circumstances should the technical properties of AI systems be employed to determine the level of regulatory scrutiny. Third, regulators should examine the entire risk chain of AI applications and strengthen structural constraints that could serve as more effective barriers to misuse.
@@
The challenge of regulating AI exemplifies many of the core sources of contention and difficulties in tech policy. Central to this discourse is the distinction between tech-neutral and tech-specific regulatory approaches. Tech-neutral regulation governs activities and consequences, while tech-specific regulation governs specific technical implementations.
The case for tech-neutral regulation has taken on new significance in the context of AI. In task-specific AI, model development is oriented to specific use cases: this task specificity facilitates compliance with tech-neutral regulation, as we argue in Section 3.1. However, for GPAI, this specificity is lost due to the separation between development and deployment. Instead, the general-purpose abilities of GPAI have led to calls for regulation at the level of the model itself, leading to shallow approaches based on model parameters and other technical proxies. Bommasani (2023) has identified two types of proxies commonly used in AI policy: (1) resource expenditure (i.e., the amount of compute in FLOPs and the amount of data in modality-specific units), and (2) model properties (i.e. capabilities benchmarks and safety benchmarks).
@@
AI regulation that relies on technical proxies risks not just becoming outdated but being both under- and over-inclusive (Crootof and Ard, 2021). [^6] Considered more precisely, this is a point about legal rules and standards, not tech neutrality (Crootof and Ard, 2021, p. 406). Where regulations are more rule-like, they can be over- and under-inclusive. We make the point here about tech-specificity because we think it is likely that tech-specificity leads to rules more often than standards. This is both because the tendency for tech-specific regulation is already motivated by a desire for easier adminsitrability and because technical standards and specifications are often stated in quantitative terms. Examples of tech-specific standard-based regulation may become more prevalent if governments, for example, mandate that LLMs have output filters that accomplish some goal like fairness or a more complicated set of goals in the vein of Anthropic’s attempt at “Constitutional AI” (Bai et al., 2022), but for now, we think the assumption holds. Regulation is under-inclusive when it leaves out systems that achieve similar concerning outcomes through alternative methods. For instance, relying on compute thresholds to identify frontier AI models may inadvertently exclude systems that cause harm through less resource-intensive methods (e.g., misinformation bots) (Newsom, 2024). Regulation is over-inclusive when it captures more systems than intended. For example, the same compute thresholds could sweep in benign applications like climate modeling that happen to share high computational demands. The challenge in defining the regulatory scope is compounded by the broad range of contexts in which the same AI systems can be deployed to serve vastly different purposes (Harris, 2021). The generality and adaptability of GPAI therefore underscore the need for regulatory frameworks that account for the specific contexts and outcomes of their use (Murdick et al., 2020).
@@
One example that highlights the limits of tech-specific regulation is the European Union’s legislation on genetically modified organisms (GMO) (Parliament and the Council of the European Union, 2001). The legislation targets crops that have been altered by recombinant DNA technology to address public concerns over potential health risks. In response, seed producers have turned away from recombinant DNA technology toward mutagenesis, using mutagenic chemicals or radiation treatments to induce random genetic changes in plants (Kaskey, 2013a). While these methods are not considered recombinant DNA technology, they achieve the same outcome of producing genetically modified crops with commercially desirable traits. Ironically, mutagenic crops may pose comparable or even greater risks than GMO crops on food safety and the environment (Kaskey, 2013b; Morris, 2007). This case demonstrates how tech-specific regulation can incentivize the development of what Dan Burk calls “perverse innovations” that “ingeniously dodge the intended outcome of regulation, while still formally adhering to the text of the regulation” (Burk, 2016).
4.2. Identifying regulatory targets by technical specifications can still make sense as a last resort.
In recent years, the AI regulatory landscape has seen a chorus of concerns over AI safety and security risks (Tobin, 2024; Lovely, 2024; Tufekci, 2025; Future of Life Institute, 2023). These concerns have led to a surge in tech-specific proposals that aim to regulate AI systems, with model properties as technical proxies for their capacity to precipitate harm in society (European Parliament, 2024; U.S. House Committee on Foreign Affairs, 2024; The White House, 2023; California Senate, 2024).
@@
However, as we highlighted in the prior section, tech-specific regulation is not without its limitations. Its application should be carefully constrained to avoid unintended consequences, such as overregulating benign use cases and creating regulatory loopholes. For this reason, we recommend that technical proxies should be employed for regulation only in narrow circumstances where three conditions are met: (1) tech-neutral regulation is demonstrably impractical in addressing the risk of concern, (2) the risk being addressed is severe, plausible, and attributable to the existence and configuration of the technology, and (3) the benefits from tech-specific regulation outweigh the potential harms. We now expand on each of these three conditions.